# Measurement of Radon-222 concentration in N2 using an activated charcoal trap

### v1

@@ 1 Introduction
The purpose of this work is to introduce a method for measuring Radon-222 ( [^2] Rn) concentration in the nitrogen (N2) cover gas systems of experiments hosted at SNOLAB, an underground science facility located 2 km deep in the Vale Creighton mine near Sudbury, Ontario. Experiments such as SNO+ [1], DEAP-3600 [2], and high-purity germanium detectors [3] use boil-off N2 from SNOLAB’s liquid nitrogen (LN) plants [4] and Grade 5 N2 cylinder bottles [5] to control the background induced by  [^2] Rn progeny. Radon-222 progeny can produce some of the most troublesome radioactive backgrounds for these experiments.  [^2] Rn is a radioactive noble gas with a half-life of 3.8 days. It is the progeny of 226Ra in the 238U decay chain as shown in Figure 1.  [^2] Rn can enter detectors through leaks, surface emanation, and diffusion.
@@
Radon-222 assay refers to the measurement of the concentration of  [^2] Rn in a given gas or water sample as well as the radon emanation and diffusion rate from a material surface. The SNO collaboration developed a technique to perform material and ultra pure water radon assays [6][7]. In this technique, radon atoms are extracted from a sample material and from water droplets, then trapped in a radon trap which is at -196 °$\degree$°C. The radon trap in these two applications is filled with bronze wool and is kept under vacuum. The bronze wool is used to stop the radon atoms by increasing the surface area, while allowing the other gasses such as N2 and O2 to pass through. The gas flow rate for these applications is $<0.1$  standard litter per minute (SLPM). This technique with the bronze wool as radon trapping mechanism initially adopted by the authors to perform N2 gas assays, in which the gas flow rate is $\geq$ 1 SLPM. However, it was observed that the radon trapping efficiency decreases with the assay duration [8]. This limits the maximum volume of N2 that can be sampled with this technique and, consequently, restricts the assay sensitivity. We hypothesize that this is due to the increasing temperature of the trap as more N2 passes through, which leads to the desorption of the radon atoms. This observation motivated us to develop and add a new radon trap filled with a material which has higher heat capacity compare to the bronze wool and use it as a pre-trap to the existing system.
@@
For this paper Calgon OVC 4 $\times$[11], which is a granular activated carbon made from coconut shell is used. The charcoal was purified at SNOLAB and then was used to make a radon trap. The background and radon trapping efficiency of the trap were measured. The latter was performed under various cryogenic temperatures, gas flow rates and flow times. To measure the radon trapping efficiency, a radon calibration source was developed and characterized. In addition, the performance of the trap was compared with that of bronze wool using underground lab air and N2 from a bottle. The outline of the paper is as follows. the experimental apparatus is detailed in Section 2. The activated charcoal radon assay and analysis methodology, as well as, the radon calibration source characterization are described in Section 3. The background and efficiency of the charcoal trap, as well as a comparison of its performance for underground lab air and N2 assays to that of bronze wool, are presented in Section 4. In addition, the performance and sensitivity of our radon system with the addition of the charcoal trap was compared with the original system without the charcoal trap. The conclusion and discussion are provided in Section 5.
@@ 2.1 Radon board
The radon boards can be used for both radon emanation measurement and gas assays. When used for gas assay the board is connected to a sample gas port thorough a vacuum fitting, while the valve to the emanation chamber is kept closed. For N2 gas flow of 1 SLPM the efficiency of the bronze wool varies from 85% to less than 3%, depending on the duration that gas flows through the trap [8]. This indicates that a 3/8” diameter tube filled with bronze wool is not a good design for performing high sensitivity gas assays.
@@ 2.2 Lucas cell counting system
Figure 4: Picture of a Lucas Cell (LC)
@@ 2.3 Radon calibration source
To measure the efficiency of the radon traps a radon calibration source was made from Nora Xp 5319 rubber floor tile [17]. We used $69\pm 2~{}\mathrm{g}$ of this material. It was cut into fourteen strips of $\mathrm{1~{}cm\times 10~{}cm}$ and was inserted in a 2.75 inch ConFlat (CF) tee. The picture of the can before and after closing the cap of the CF flange is shown in Figure 6. The calibration source was later characterized by the SNOLAB surface radon board. The radon emanation results of this source are presented in Section 3.5. The apparatus to connect the calibration source to the radon board is shown in Figure 3 c.
@@ 3.1 Activated charcoal trap testing
The efficiency of  [^2] Rn adsorption and desorption of the activated charcoal trap is measured across a range of flow rates from 1 SLPM to 7 SLPM. A combination of  [^2] Rn emanation from the test source (calibration can) and N2 as a carrier gas is used to mimic underground cover gas systems. In this case, the first transfer is through the activated charcoal trap using the vacuum pump while the trap is cooled. Initially, LN was used in order to cool down the trap, but it was discovered that LN slows the N2 molecules, resulting in high pressure ($>$30 PSI) and subsequently increased uncertainty on the assay results. To overcome this issue, alcohol slush — a mixture of reagent alcohol and LN — was used as the cryo-liquid. The temperature of the cryo-liquid was varied in order to measure the efficiency of trap versus temperature. The temperature of the external wall of the U-tube was measured using a dedicated sensor. The subsequent transfers are the same as described earlier (Charcoal $\rightarrow$ Trap A $\rightarrow$ Trap B $\rightarrow$ LC).
@@ 3.3 Radon board efficiency measurement
The efficiency of the radon board is combination of the Trap A efficiency, cryo-pumping transfer efficiency from Trap A to Trap B, and transfer efficiency from Trap B to the LC. As the transfer of radon atoms from Trap B to LC is done through volume sharing, the latter can be calculated by looking at the value of the pressure transducer (PB) before and after opening the valve between the LC and Trap B. The transfer efficiency from Trap B to LC is calculated to be $83\pm 1$% and $81\pm 1$% for the surface and UG boards, respectively. The uncertainty on this measurement was found by performing five measurements for each board. To measure combined transfer efficiency, an LC was filled with radon atoms and then counted using the counting system to determine the exact amount of radon atoms in the cell. The LC then was then attached to Vsample via a quick connect. While Trap A was at LN temperature and under vacuum, $V_{sample}$ was opened and the LC was pumped through Trap A for an hour, following a procedure similar to that used for performing radon assays. By the end of the hour the LC pressure reached the ultimate pressure of the vacuum pump ($5.0\times 10^{-3}$ mbar). The rest of the procedure was followed to direct the radon atoms to Trap B. In order to remove any systematic uncertainty related to LC efficiency, the LC was detached from Vsample and was attached to the LC port to transfer radon atoms from Trap B to the same cell. The total efficiency of the board is defined as:
@@
where $\epsilon_{vol}$ is Trap B to LC radon transfer efficiency due to volume sharing. This gives the combination of Trap A and Trap B efficiency in trapping radon atoms which is $96^{+4}_{-8}\%$ and $94^{+6}_{-7}\%$.
@@ 4.1 Background measurement of the charcoal trap
Two sets of measurements were performed in order to determine the charcoal trap  [^2] Rn background. For the first set, the charcoal trap was purged with N2 and was baked for 20 minutes. It was left to be emanated under vacuum in room temperature for 4 and 17 days. Then the charcoal trap was connected to $V_{sample}$ and radon atoms in the charcoal were transferred to Trap A using the vacuum pump for an hour. The charcoal trap was heated to $150~{}\degree$C during the transfer, to increase the thermal energy of the radon atoms to break the Van der Waals bonds with charcoal molecules. The transfer of radon atoms to Trap B and an LC was performed using same procedure described in the methodology. The two emanation rate measurements were $36\pm 6$ radon atoms/day and $34\pm 6$ radon atoms/day. The error bars consist of a 14% systematic uncertainty, as well as the statistical uncertainty, which was added in quadrature.
The second set of measurements was performed in order to determine the number of radon atoms in charcoal in the same operating conditions under which gas assays are performed. For normal assays this trap is used in total for two hours. During the first hour the trap is under vacuum and at a cryogenic temperature of -70 °$\degree$°C. In the second hour, the trap is warmed up to $150~{}\degree$C . For this work a cleaning procedure was developed in to remove any residual radon atoms from previous measurements. Initially, the trap was purged with N2 from a bottle, but later it was discovered that N2 purging elevates the background of the trap, indicating that N2 from a bottle is itself is a source of radon. N2 purging was therefore removed from the cleaning procedure. It was also determined that in order to remove all residual radon atoms remaining in charcoal pores from previous assays the trap needs to be baked for at least two hours at $150~{}\degree$C. The background of the charcoal trap was then measured as it was under vacuum and was cooled to $-70\degree$ C for one hour, then later heated at 150 °$\degree$°C for one hour. This is similar to the operating conditions of this trap. This measurement was performed twice and the number of radon atoms was found to be $4.6\pm 2.8$, using the average value of the two measurements. This background level is used for the all charcoal trap assays reported in this paper. SNOLAB’s Well detector set the limit of $<$ 33 mBq/kg [14], corresponding to $<$ 2.6 radon atoms per hour of assay time for the 22 g charcoal which is in the U-trap. The results between two methods are less than two sigma away from each other. For the Well detector’s measurement, 3.2 g of charcoal was used in comparison with 22 g used in the trap. Therefore the results are consistent statistically.
@@ 4.2.1 Efficiency versus temperature
Charcoal trap efficiency was measured from a temperature of -196 °$\degree$°C (LN temperature) to -20°$\degree$°C for a mass flow rate of 5 SLPM and one hour of assay time. The results are shown in Figure 8. The variation of the temperature shown on the x-axis of the plot is due to the increasing of temperature during the one hour of assay. This was due to the fact that the alcohol slush was made before starting the assays and the top of the coolant used was exposed to room temperature during the procedure. The overlapping bins in the graph is due to the overlapping temperatures. Therefore the measurements in the overlapping regions are not statistically independent. From this graph it was concluded that a cryo-liquid temperature of $<$$-50~{}\degree$C yields a 100% trapping efficiency. It was decided that the trap should be operated at a starting alcohol slush temperature of $-75~{}\degree$C and the temperature of the cryo-liquid is monitored during an assay to ensure the temperature is maintained to be $<-50~{}\degree$C.
@@ 4.2.2 Efficiency versus mass flow rate and assay duration
Tests were performed using a one hour assay duration for flow rate of 1 SLPM, 3 SLPM, 5 SLPM and 7 SLPM while the slush temperature was kept between $-80~{}\degree$C to $-65~{}\degree$C. It was found that the relative efficiency of the charcoal trap stayed constant for these flows and was consistent with 100%.
Maintaining the temperature of the cryo-liquid in the region $<$$-50~{}\degree$C for an assay duration of longer than two hours was challenging and therefore it was decided to set the maximum assay time to two hours. It was found out that for a flow rate of 5 L/min the efficiency of charcoal trap for a two hour assay was $(87.5\pm 12.0)\%$.
@@ 4.4 N2 gas assay sensitivity using charcoal trap
4.5 Measuring radon level in $N_{2}$ bottle using charcoal and bronze trap
We measured the concentration of radon in a Grade 5 N2 bottle [5] using the bronze wool and activated charcoal trap, following the procedures described above. For the bronze wool measurement, we performed a 40 L gas assay at flow rate of 1 SLPM. For the charcoal trap measurement, we extracted 300 L at a flow rate of 5 SLPM. The results obtained with the bronze trap were consistent with LC background and corresponded to $2.3\pm 4.2$ Rn atoms/L of $N_{2}$. The results using charcoal trap was $0.6\pm 0.1$ Rn atoms/L of N2. N2 assay results confirms that we have improved our capability at SNOLAB to measure low concentration of  [^2] Rn in N2.

### v2

@@ 1 Introduction
The purpose of this work is to introduce a method for measuring Radon-222 ( [^2] Rn) concentration in the nitrogen (N2) cover gas systems of experiments hosted at SNOLAB, an underground science facility located 2 km deep in the Vale Creighton mine near Sudbury, Ontario. Experiments such as SNO+ [1], DEAP-3600 [2], and high-purity germanium detectors [3] use boil-off N2 from SNOLAB’s liquid nitrogen (LN) plants [4] and Grade 5 N2 cylinder cylinders [5] to control the background induced by  [^2] Rn progeny. Radon-222 progeny can produce some of the most troublesome radioactive backgrounds for these experiments.  [^2] Rn is a radioactive noble gas with a half-life of 3.8 days. It is the progeny of 226Ra in the 238U decay chain as shown in Figure 1.  [^2] Rn can enter detectors through leaks, surface emanation, and diffusion.
@@
Radon-222 assay refers to the measurement of the concentration of  [^2] Rn in a given gas or water sample as well as the radon emanation and diffusion rate from a material surface. The SNO collaboration developed a technique to perform material and ultra pure water radon assays [6][7]. In this technique, radon atoms are extracted from a sample material or from water droplets, then trapped in a radon trap which is at -196 °$\degree$°C. The radon trap in these two applications is filled with bronze wool and is kept under vacuum. The bronze wool is used to stop the radon atoms by increasing the surface area, while allowing the other gasses such as N2 and O2 to pass through. The gas flow rate for these applications is $<0.1$  standard litter per minute (SLPM). This technique with the bronze wool as radon trapping mechanism initially adopted by the authors to perform N2 gas assays, in which the gas flow rate is $\geq$ 1 SLPM. However, it was observed that the radon trapping efficiency decreases with the duration that gas flows through the trap[8]. This limits the maximum volume of N2 that can be sampled with this technique and, consequently, restricts the assay sensitivity. We hypothesize that this is due to the increasing temperature of the trap as more N2 passes through, which leads to the releasing of the radon atoms. This observation motivated us to develop and add a new radon trap filled with a material which has higher heat capacity compared to bronze wool and use it as a pre-trap to the existing system.
@@
For this paper Calgon OVC 4 $\times$[11], which is a granular activated carbon made from coconut shell is used. The charcoal was purified at SNOLAB and then was used to make a radon trap. The background and radon trapping efficiency of the trap were measured. The latter was performed under various cryogenic temperatures, gas flow rates and flow times. To measure the radon trapping efficiency, a radon calibration source was developed and characterized. In addition, the performance of the trap was compared with that of bronze wool using underground lab air and N2 from a cylinder. The outline of the paper is as follows. the experimental apparatus is detailed in Section 2. The activated charcoal radon assay and analysis methodology, as well as, the radon calibration source characterization are described in Section 3. The background and efficiency of the charcoal trap, as well as a comparison of its performance with underground lab air and N2 assays to that of bronze wool, are presented in Section 4. In addition, the performance and sensitivity of our radon system with the addition of the charcoal trap was compared with the original system without the charcoal trap. The conclusion and discussion are provided in Section 5.
@@ 2.1 Radon board
The radon boards can be used for both radon emanation measurement and gas assays. When used for gas assay the board is connected to a sample gas port through a vacuum fitting, while the valve to the emanation chamber is kept closed. For N2 gas flow of 1 SLPM the efficiency of the bronze wool varies from 85% to less than 3%, depending on the duration that gas flows through the trap [8]. This indicates that a 3/8” diameter tube filled with bronze wool is not a good design for performing high sensitivity gas assays.
@@ 2.2 Lucas cell counting system
Figure 4: Picture of a Lucas Cell (LC) and one of the PMTs used for counting.
@@ 2.3 Radon calibration source
To measure the efficiency of the radon traps a radon calibration source was made from Nora Xp 5319 rubber floor tile [17]. We used $69\pm 2~{}\mathrm{g}$ of this material. It was cut into fourteen strips of $\mathrm{1~{}cm\times 10~{}cm}$ and was inserted in a 2.75 inch ConFlat (CF) tee. The picture of the can before and after closing the cap of the CF flange is shown in Figure 6. The calibration source was later characterized using the SNOLAB surface radon board. The radon emanation results of this source are presented in Section 3.5. The apparatus to connect the calibration source to the radon board is shown in Figure 3c.
@@ 3.1 Activated charcoal trap testing
The efficiency of  [^2] Rn adsorption and desorption of the activated charcoal trap is measured across a range of flow rates from 1 SLPM to 7 SLPM. A combination of  [^2] Rn emanation from the test source (calibration can) and N2 as a carrier gas is used to mimic underground cover gas systems. In this case, the first transfer is through the activated charcoal trap using the vacuum pump while the trap is cooled. Initially, LN was used in order to cool down the trap, but it was discovered that LN slows the N2 molecules, resulting in a high pressure within the trap ($>$30 PSI) and subsequently increased uncertainty on the assay results. To overcome this issue, alcohol slush — a mixture of reagent alcohol and LN — was used as the cryo-liquid. The temperature of the cryo-liquid was varied in order to measure the efficiency of trap versus temperature. The temperature of the external wall of the U-tube was measured using a dedicated sensor. The subsequent transfers are the same as described earlier (Charcoal $\rightarrow$ Trap A $\rightarrow$ Trap B $\rightarrow$ LC).
@@ 3.3 Radon board efficiency measurement
The efficiency of the radon board is the combination of Trap A efficiency, cryo-pumping transfer efficiency from Trap A to Trap B, and transfer efficiency from Trap B to the LC. As the transfer of radon atoms from Trap B to LC is done through volume sharing, the latter can be calculated by looking at the value of the pressure transducer (PB) before and after opening the valve between the LC and Trap B. The transfer efficiency from Trap B to LC is calculated to be $83\pm 1$% and $81\pm 1$% for the surface and UG boards, respectively. The uncertainty on this measurement was found by performing five measurements for each board. To measure combined transfer efficiency, an LC was filled with radon atoms and then counted using the counting system to determine the exact amount of radon atoms in the cell. The LC was then attached to Vsample via a quick connect. While Trap A was at LN temperature and under vacuum, $V_{sample}$ was opened and the LC was pumped through Trap A for an hour, following a procedure similar to that used for performing radon assays. By the end of the hour long transfer, the LC pressure reached the ultimate pressure of the vacuum pump ($5.0\times 10^{-3}$ mbar). The rest of the procedure was followed to direct the radon atoms to Trap B. In order to remove any systematic uncertainty related to LC efficiency, the LC was detached from Vsample and was attached to the LC port to transfer radon atoms from Trap B to the same cell. The total efficiency of the board is defined as:
@@
where $\epsilon_{vol}$ is Trap B to LC radon transfer efficiency due to volume sharing. This gives the combination of Trap A and Trap B efficiency in trapping radon atoms to be $96^{+4}_{-8}\%$ and $94^{+6}_{-7}\%$ for surface and underground board, respectively.
@@ 4.1 Background measurement of the charcoal trap
Two sets of measurements were performed in order to determine the charcoal trap  [^2] Rn background. For the first set, the charcoal trap was purged with N2 and was baked for 20 minutes. It was left to be emanated under vacuum at room temperature for 4 and 17 days. Then the charcoal trap was connected to $V_{sample}$ and emanated radon atoms from the charcoal were transferred to Trap A using the vacuum pump for an hour. The charcoal trap was heated to $150~{}\degree$C during the transfer, to increase the thermal energy of the radon atoms to break the Van der Waals bonds with charcoal molecules. The transfer of radon atoms to Trap B and an LC was performed using the same procedure described in the methodology. The two emanation rate measurements were $36\pm 6$ radon atoms/day and $34\pm 6$ radon atoms/day. The error bars consist of a 14% systematic uncertainty, as well as the statistical uncertainty, which was added in quadrature.
The second set of measurements was performed in order to determine the number of radon atoms in charcoal in the same operating conditions under which gas assays are performed. For normal assays this trap is used in total for two hours. During the first hour the trap is under vacuum and at a cryogenic temperature of -70 °$\degree$°C. In the second hour, the trap is warmed up to $150~{}\degree$C . For this work a cleaning procedure was developed to remove any residual radon atoms from previous assays. Initially, the trap was purged with N2 from a cylinder, but later it was discovered that N2 purging elevates the background of the trap, indicating that N2 from a cylinder is itself is a source of radon. N2 purging was therefore removed from the cleaning procedure. It was also determined that in order to remove all residual radon atoms remaining in charcoal pores from previous assays the trap needs to be baked for at least two hours at $150~{}\degree$C. The background of the charcoal trap was then measured as it was under vacuum and was cooled to $-70\degree$ C for one hour, then later heated at 150 °$\degree$°C for one hour. This is similar to the operating conditions of this trap. This measurement was performed twice and the number of radon atoms was found to be $4.6\pm 2.8$, using the average value of the two measurements. This background level is used for the all charcoal trap assays reported in this paper. SNOLAB’s Well detector set the limit of $<$ 33 mBq/kg [14], corresponding to $<$ 2.6 radon atoms per hour of assay time for the 22 g charcoal which is in the U-trap. The results between two methods are less than two sigma away from each other. For the Well detector’s measurement, 3.2 g of charcoal was used in comparison with 22 g used in the trap. Therefore the results are consistent statistically.
@@ 4.2.1 Efficiency versus temperature
Charcoal trap efficiency was measured from a temperature of -196 °$\degree$°C (LN temperature) to -20°$\degree$°C for a mass flow rate of 5 SLPM and one hour of assay time. The results are shown in Figure 8. The variation of the temperature for each data points (green bins) is due to the increasing of temperature of the alcohol slush during the one hour of assay. This was due to the fact that the alcohol slush was made before starting the assays and the top of the coolant used was exposed to room temperature during the procedure. The overlapping bins in the graph is due to the overlapping temperature of alcohol slush. Therefore the measurements in the overlapping regions are not statistically independent. From this graph it was concluded that the alcohol slush temperature of $<$$-50~{}\degree$C yields a 100% trapping efficiency. It was decided that the trap should be operated at a starting alcohol slush temperature of $-80~{}\degree$C and the temperature of the alcohol slush is monitored during an assay to ensure the temperature is maintained to be $<-50~{}\degree$C.
@@ 4.2.2 Efficiency versus mass flow rate and assay duration
Tests were performed using a one hour assay duration for flow rate of 1, 3, 5 and 7 SLPM while the slush temperature was kept between $-80~{}\degree$C to $-65~{}\degree$C. It was found that the relative efficiency of the charcoal trap stayed constant for these flows and was consistent with 100%.
Maintaining the temperature of the alcohol slush in the region $<$$-50~{}\degree$C for an assay duration of longer than two hours was challenging and therefore it was decided to set the maximum assay time to two hours. It was found out that for a flow rate of 5 L/min the efficiency of charcoal trap for a two hour assay was $(87.5\pm 12.0)\%$.
@@ 4.4 N2 gas assay sensitivity using charcoal trap
4.5 Measuring radon level in $N_{2}$ cylinder using charcoal and bronze trap
We measured the concentration of radon in a Grade 5 N2 cylinder [5] using the bronze wool and activated charcoal trap, following the procedures described above. For the bronze wool measurement, we performed a 40 L gas assay at flow rate of 1 SLPM. For the charcoal trap measurement, we extracted 300 L at a flow rate of 5 SLPM. The results obtained with the bronze trap were consistent with LC background and corresponded to $2.3\pm 4.2$ Rn atoms/L of $N_{2}$. The results using charcoal trap was $0.6\pm 0.1$ Rn atoms/L of N2. N2 assay results confirms that we have improved our capability at SNOLAB to measure low concentration of  [^2] Rn in N2.